What is the total rent cost of all the machines used by UniVEND each month? Report the total monthly fixed cost basis. How many products have to be sold each month to cover the monthly fixed cost basis? b) Price Average Earning Turkish Corporate Tax Rate 20.0% average price 2.00 T per product Variable Costs Average Cost Average Sales Number 1.65 TL 0.05 TL per product per product averoge soles (monthly) 7,500.00 per machine average cost variable overnead Fixed Costs Cost Vending Machines Necta Snaccky Max Necta Snaccky Elektral Midi Bufe Elektral Maxi Bufe Vendo Cold Drink Number of Machines Vending Machines Rent Necta Snaccky Max Necta Snaccky Elektral Midi Bufe Elektral Maxi Bufe Vendo Cold Drink 250.00 TL 250.00 TL 350.00 TL 650.00 TL 450.00 TL 200.00 TL 4,000.00 TL monthly -per machine monthly -per machine monthly - per machine monthly -per machine monthly- per machine per machine for all machines 13 Utilities (monthly) Fixed Overhead(monthly) Selling & Admin Costs Cost Monagement Salanes Selling and Administration 3,300.00 TL 18,000.00 TL monthly annually
Answer #1

1a)Total rent costs=(no of machines*per machine rent)
=(5*250)+(3*250)+(5*350)+(3*650)+(13*450)
=11550
b)total monthly fixed costs=11550+(200*(5+3+5+3+13))+4000+3300+(18000/12)
=26150
c)No of products needs to be sold=fixed costs/(sales-variable costs)
=26150/(2-(1.65+0.05))
=87167(rounded)